# Precession and higher order multipoles in binary black holes (and lack thereof)

### v1

@@ III.1 Evidence for precession and higher order multipole moments
Figure 1 shows the cumulative distribution of p-values for precession and higher order multipoles for all confident binary black hole mergers observed by the LVK. The dashed black line shows the expected uniform distribution of p-values under the null hypothesis that there is no signal from precession or higher order multipoles, i.e. under the assumption that Gaussian noise alone produced any precession and higher order multipole signatures in the observed GWs. We also show the expected $1\sigma$, $2\sigma$ and $3\sigma$ scatter in p-values under the null hypothesis for a population of 69 binary black holes. The blue and orange curves show the actual distribution of p-values for precession and higher order multipoles, respectively.
@@
In Figure 1, we see a small excess of low- significance events that have more significant p-values for precession than expected. This trend was also seen in Ref.[54]. Although Ref.[54] noted some possible reasons for this, we expect that this small excess is a statistical fluctuation. This may also be the reason why other studies infer precession in the binary black hole population, despite only one candidate with significant evidence [15].
Only GW190814 and GW190412 show strong evidence for higher order multipoles, specifically the $(\ell,m)=(3,3)$ multipole, with probabilities of these events coming from sources with no higher order multipoles of $10^{-11}$ and $2\times 10^{-4}$, respectively. Based on the expected distribution of p-values from a population of 69 binary black holes with no precession or higher order multipoles, these are a $>5\sigma$ and just over a $4\sigma$ detection, respectively. This supports the conclusions found in Refs. [51, 38, 53, 52]. The strong evidence for higher order multipole moments indicates that both binaries have unequal component masses; indeed a full Bayesian analysis infers mass ratios $\sim 9:1$ and $\sim 3.5:1$ respectively [51, 38]. The strong evidence for higher order multipoles in GW190412 has also allowed for measurements of the sources recoil velocity [86].
Figure 2: Posterior distributions for the precession phase, $\hat{\phi}_{JL}$ [87], for the observed binary black hole mergers in GWTC-3, coloured by the p-value for the evidence of precession $P_{p}$. As in Ref. [87], we artificially centre each posterior distribution by subtracting the circular mean.
GW190929_012149 is the next most significant candidate for an observation of higher order multipoles with a probability of no higher order multipole content in the underlying source of $0.02$, i.e. 1 in 50 chance that Gaussian noise is responsible for the observed higher order multipole effects. Although this was deemed marginal in Ref. [54], the larger population presented here means that this is now in line with expectations. Marginal evidence for higher order multipoles was previously reported in GW151226 [55], GW170729 [1, 56, 57] and GW190519_153544 [54]. Our analysis finds that these events have p-values $0.3$, $0.1$ and $0.05$ respectively. When comparing to a population of 69 events, we see no statistical evidence for higher order multipoles in GW151226, GW170729 and GW190519_153544.
@@ III.2 Correlation between the precession phase and probability of precession
The precession phase, $\phi_{JL}$ [^2] The precession phase, $\phi_{JL}$, is sometimes denoted in the literature as $\alpha$. We denote the $2\pi$-periodic version of the precession phase by $\hat{\phi}_{JL}$ to match the notation used in Ref. [87]., tracks the azimuthal angle of the orbital angular momentum $\mathbf{L}$ around the total angular momentum $\mathbf{J}$. As described in Ref. [70], $\phi_{JL}$ imparts an overall phase shift to each of the waveform components in the harmonic decomposition. When precession is not observable, the precession phase cannot be constrained and Bayesian analyses will return an uninformative measurement of $\phi_{JL}$. However, if the GW has observable precession and no higher order multipoles, Equation 1 reduces to $h\approx h^{\prime}+h_{p}$, and the precession phase can be measured independently from the orbital phase. Since the orbital phase remains the same for both waveform components, while the contribution from the precession phase differs, constraints can be placed on the value of $\phi_{JL}$.
Recently Ref. [87] found that the precession phase, or rather a $2\pi$-periodic version of the precession phase $\hat{\phi}_{JL}$, is recovered surprisingly well for several candidates, most notably GW200129, GW190412 and GW151226. In Figure 2, we reproduce the analysis of Ref. [87] but look for correlations between the precession p-value and the precision of the measurement of $\hat{\phi}_{JL}$.
As expected we see that, in general, a smaller precession p-value correlates with a tighter constraint on $\hat{\phi}_{JL}$. GW200129, GW190412 and GW151226 have the tightest constraints on $\hat{\phi}_{JL}$, and they have small precession p-values: $4\times 10^{-5}$, $0.04$ and $0.12$. We see one noticeable exception though: GW191109_010717. This candidate has a p-value of 0.018 with a relatively poorly constrained $\hat{\phi}_{JL}$.
@@ III.3 Comparing observations with expectations from population estimates
We now turn to comparing the fraction of observed events that display signatures of precession or higher order multipole moments with expectations from the population model described below. We check for consistency by sampling binaries from the population model, calculating the expected distributions of p-values for precession and higher order multipole moments, and comparing these distributions with observations. In this work we assume a Powerlaw+peak mass model [88], a Default spin model [89] and a Powerlaw redshift model [90], all described in Appendix B of Ref. [15]. We randomly draw 1000 samples from the posterior distribution on the hierarchical model parameters obtained in Ref. [15]; in effect we simulate 1000 different Universes, each consistent with the data. We then randomly sample merging binaries from each Universe. We hereafter refer to this population model as the LVK population model.
@@
Not all binaries drawn from population models are observable with current template-based search pipelines adopted by the LVK [91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103]. Current template-based search pipelines adopted by the LVK use templates which restrict spins to be aligned with the orbital angular momentum and neglect the presence of higher order multipole moments (although see Refs.[81, 104, 105, 106, 12] for possible solutions for template-based search pipelines, Refs. [107, 108, 109] for search pipelines that do not use templates, and Ref. [110] which demonstrates that non-template based search pipelines may be more sensitive to precessing binaries than current template-based search pipelines). This means that current template-based searches are not equally sensitive to all binaries, and binaries with significant evidence for precession and higher order multipole moments may remain undetected.
@@
As discussed in Section II, multiple Bayesian inference analyses are needed to calculate p-values for precession and higher order multipole moments for a single randomly chosen binary. Since performing Bayesian inference is a computationally expensive task, let alone analysing each binary with multiple models, we only analyse randomly chosen binaries with $\rho_{\mathrm{p}}>1$ and $\rho_{33}>1$. Although we expect similarities between populations with $\rho_{\mathrm{p}}>1$ and $\rho_{33}>1$, for example asymmetric component masses [42, 50], we treat them as separate populations in this work. This means that we separately draw two populations: one where binaries have $\rho_{\mathrm{p}}>1$ and another where binaries have $\rho_{33}>1$. We find that $11.5\%$ and $5.6\%$ of the detectable binaries sampled from the LVK population model have $\rho_{\mathrm{p}}>1$ and $\rho_{33}>1$, respectively. To ensure a consistent number of binaries in each population, we initially draw $100$ detectable binaries from each Universe in the LVK population model, giving a total of $10^{5}$ randomly chosen binaries, and downsample to 1000 and 2000 binaries for the precession and higher order multipole analyses respectively. We therefore estimate the expected distribution of p-values for precession and higher order multipole moments from population models based on $\sim 120$ simulated GW signals. We take care to renormalize the p-value distributions, based on the expected number of binaries with $\rho_{\mathrm{p}}>1$ and $\rho_{33}>1$.
@@
Figure 3 compares the cumulative distribution of p-values for precession and higher order multipole moments with random draws from population estimates. The current LVK population model excludes GW190814 [15], so we also excluded this event from the cumulative distribution. We see that the LVK population model accurately predicts the number of observed events with significant evidence for precession and higher order multipole moments, with the distribution of observed p-values lying comfortably within the $1\sigma$ scatter from the population samples. We see large error bars associated with the expectation from population estimates, with the most significant binaries having a large spread of p-values: $10^{-8}<P_{\mathrm{p}}<4\times 10^{-3}$ for precession and $10^{-11}<P_{33}<10^{-2}$ for the $(\ell m)=(3,3)$ multipole, with ranges quoted at the $1\sigma$ level. This is a consequence of analysing a limited number of randomly chosen binaries owing to computational cost. Unsurprisingly, when we include GW190814 in the cumulative distribution, the observed distribution now lies outside of the $1\sigma$ uncertainty. This is expected since current population estimates have vanishingly small probabilities for drawing binaries with extreme mass ratio configurations.
@@
In Figure 4 we show the posterior distribution for the inferred primary spin magnitude, $a_{1}$, and mass ratio, $q=m_{2}/m_{1}\leq 1$, for a binary chosen from the population model. This binary was chosen as it was inferred to have a significant precession p-value: $P_{\mathrm{p}}<4\times 10^{-5}$, and hence a more significant p-value than GW200129. We see that the inferred posterior distribution resembles GW200129, with near extreme spin and $q\sim 0.5$; GW200129 was inferred to have $a_{1}=0.9^{+0.1}_{-0.5}$ and $q=0.6^{+0.4}_{-0.2}$ [43]. Figure 4 also shows the posterior distribution for another randomly chosen binary from the population model; this binary was chosen as it was inferred to have a significant higher order multipole p-value: $P_{\mathrm{33}}<2\times 10^{-4}$, and hence a more significant p-value than GW190412. Here, we see that the inferred inclination angle, $\theta_{JN}$, and mass ratio, $q$, resembles GW190412; GW190412 was inferred to have $\theta_{JN}=0.71^{+0.31}_{-0.24}$ and $q=0.28^{+0.12}_{-0.07}$ [38]. In both cases, the true values lie within the inferred 90% credible intervals, but there is a small bias, which is likely a consequence of the specific Gaussian noise realisation chosen. Figure 4 illustrates that GW200129-like and GW190412-like signals are possible in the LVK population model.
@@
Finally, we can use the distribution of p-values obtained from random draws from the LVK population model to calculate the probability of observing binaries with significant evidence for precession or higher order multipole moments. We find that there is a 1 in 50 chance that a detectable binary from the LVK population model will have evidence for precession that is at least as significant as for GW200129, and a 1 in 70 chance that a detectable binary will have evidence of higher order multipoles that is at least as significant as for GW190412. Both are consistent with the observation of 1 binary in 68 (removing GW190814).
@@
In this work, we chose to only compare our observations with one population model. However, we note that since we have only observed a single event with significant evidence for precession, and two events with significant evidence for higher order multipole moments, any population model that includes a subpopulation of binaries with a $\sim 2\%$ chance of detectable precession and higher order multipole moments will likely be consistent with our observations. This includes the LVK population model, but this is not the only possible model consistent with the data, and care should be taken in making strong statements about the population-wide distribution of spin-orbital misalignments and mass ratios based on a handful of events.
@@ IV Discussion
In this work we investigated whether binary black holes observed by the LVK show evidence for precession and higher order multipoles. We took advantage of the harmonic decomposition of a GW [70] and directly calculated the probability that the observed SNR in precession and higher order multipoles can be explained by Gaussian noise alone. Although the SNRs in precession and higher order multipoles have been used to help explain the physics of several GW candidates [38, 51], they are generally underutilised in population analyses.
@@
We found that only one event shows substantial evidence for precession: GW200129. Meanwhile, two events show substantial evidence for higher order multipoles: GW190412 and GW190814. This implies that any evidence for precession and higher order multipoles in other GW signals is consistent with random fluctuations caused by noise. We then compared our observations with expectations from population models to ensure consistency. We found that the current LVK population model [15] accurately predicts the number of observed events with significant evidence for precession and higher order multipole moments. We found that a binary with significant evidence for precession will occur once in every $\sim 50$ detections, and a binary with significant evidence for higher order multipole moments will occur once in every $\sim 70$ observations if sources are drawn from the LVK population model. Both fractions are consistent with observations, and the latter is consistent with forecasts from Monte Carlo simulations [113]. However, we note that since significant evidence for precession and higher order multipole moments has only been confidently observed in three events, any population model that includes a subpopulation of binaries with $\sim 2\%$ chance of detectable precession and higher order multipoles will likely be consistent with the data.
@@
Although we only find one candidate with significant evidence for precession, this does not necessarily mean that most of the observed binary black holes are non-precessing. Rather it means that the majority of the observed binary black holes do not have individually-observable precession. This could indicate a true lack of precession in binaries whose components spins are aligned with the orbital angular momentum; or component spins that are too small in magnitude for precession to be observable despite misalignment; or extrinsic parameters that are unfavourable for the detectability of precession. Consequently, all we can conclusively say about individual events is that we have observed precession in one candidate, GW200129. GW200129 is therefore likely formed through dynamic encounters (although there are cases where the isolated channel can produce binaries with high spins and significant precession [114]). A possibility is that GW200129 was formed through a hierarchical merger, where spin-induced orbital precession is likely; one component is predicted to be highly spinning and misaligned with the orbit. Indeed, signatures of hierarchical mergers may have already been found among GW observations [26, 115, 116]. Other options include triple black hole systems [117], or formation in active galactic nuclei [118]. Given that we cannot make any statements about precession of the other individual GW candidates, we are unable to comment on their formation mechanism. We note that recently Ref. [119] conducted a dedicated search for precessing binaries using data from O3, and did not find any evidence for additional GW candidates, despite an increase in sensitivity. This similarly implies that the majority of the observed binary black holes do not have individually-observable precession. We also highlight that Ref. [120] equivalently found that precession is only individually identified in GW200129 (using another observable), however they argued that at least $\sim 80\%$ of observed binaries must be precessing, even though it is not possible to identify which ones through individual observations.
@@
We only find two candidates with significant evidence for higher order multipoles, both of which have conclusively unequal component masses. Given that GW190412 does not have observable precession, the isolated formation channel is possible: although the isolated formation channel favours comparable component masses [121, 122], binaries with $q\leq 0.5$ are only disfavoured by a factor of $\sim 10$ and therefore unsurprising given the current population [123, 124, 125]. Hierarchical mergers are also possible, although we would then have expected to observe significant evidence for precession. Both isolated binary evolution models [123, 124, 126] and dynamical formation scenarios [127] struggle to reproduce the extreme mass ratio of GW190814. All binaries (excluding exactly equal mass ratio binaries) radiate power in higher order multipoles, so the lack of observable power in most of the GW candidates is partly a consequence of their extrinsic parameters.
@@
Throughout this work we only considered the observability of precession and higher order multipole moments in binary black holes observed with ground-based detectors operating at sensitivities. Previous works have investigated their observability for next generation detectors, such as Einstein Telescope [128] and Cosmic Explorer [129], and found that higher order multipoles in particular will be observable in a larger fraction of detected GW signals [130, 131, 132]. Assuming current population models and a representative redshift distribution for a next generation detector, such as the distributions proposed in Refs. [133, 134], we predict that Einstein Telescope will observe higher order multipoles once in every 20 candidates, while Cosmic Explorer will observe higher order multipoles once in every 10 candidates. On the other hand, Einstein Telescope will observe precession once in 5 candidates, while Cosmic Explorer will observe precession once in every 4 candidates. Here we assumed that a candidate will have observable precession and higher order multipoles if $\rho_{\mathrm{p}}$ and $\rho_{33}>2.1$, following the analysis in Ref. [41]. We also assumed that all candidates with SNR greater than 8 are detectable. We did not assume a network of next generation detectors, unlike Ref. [130], so these predictions are a conservative lower bound. Owing to the larger number of confident detections, we expect that next generation detectors will be influential in deciphering the formation mechanism of binary black holes.
@@
The method we presented in this paper straightforwardly and clearly identifies which GW candidates show significant evidence for precession and higher order multipole moments. This method can also be used to verify whether population estimates are consistent with observations. We therefore suggest that future population analyses make use of this approach.

### v2

@@ III.1 Evidence for precession and higher order multipole moments
Figure 1 shows the cumulative distribution of p-values for precession and HoMs for all confident binary black hole mergers observed by the LVK. The dashed black line shows the expected uniform distribution of p-values under the null hypothesis that there is no signal from precession or HoMs, i.e. under the assumption that Gaussian noise alone produced any precession and HoM signatures in the observed GWs. We also show the expected $1\sigma$, $2\sigma$ and $3\sigma$ scatter in p-values under the null hypothesis for a population of 69 binary black holes estimated from a binomial distribution, see e.g. Ref. [88] for details). The blue and orange curves show the actual distribution of p-values for precession and HoMs, respectively.
@@
In Figure 1, we see an excess of low-significance events that have smaller p-values for precession than expected. This trend was also seen in Ref. [54]. Although Ref. [54] noted some possible reasons for this, we expect that this small excess is a statistical fluctuation. In particular, we note that the Jensen-Shannon divergence is relatively flat near the minimum which means that small variations in the $\rho_{p}$ distribution can lead to large differences in the inferred p-value.
Only GW190814 and GW190412 show strong evidence for HoMs, specifically the $(\ell,m)=(3,3)$ multipole, with probabilities of these events coming from sources with no HoMs of $10^{-11}$ and $2\times 10^{-4}$, respectively. Based on the expected distribution of p-values from a population of 69 binary black holes with no HoMs, these are a $>5\sigma$ and just over a $4\sigma$ detection, respectively. This supports the conclusions found in Refs. [51, 38, 53, 52]. The strong evidence for HoMs indicates that both binaries have unequal component masses; indeed a full Bayesian analysis infers mass ratios $\sim 9:1$ and $\sim 3.5:1$ respectively [51, 38]. The strong evidence for HoMs in GW190412 has also allowed for measurements of the sources recoil velocity [93].
Figure 2: Posterior distributions for the precession phase, $\hat{\phi}_{JL}$ [94], for the observed binary black hole mergers in GWTC-3, coloured by the p-value for the evidence of precession. As in Ref. [94], we artificially centre each posterior distribution by subtracting the circular mean.
GW190929_012149 is the next most significant candidate for an observation of HoMs with a probability of no HoM content in the underlying source of $0.02$, i.e. 1 in 50 chance that Gaussian noise is responsible for the observed HoM effects. Although this was deemed marginal in Ref. [54], the larger population presented here means that this is now in line with expectations. Marginal evidence for HoMs was previously reported in GW151226 [55], GW170729 [1, 56, 57] and GW190519_153544 [54]. Our analysis finds that these events have p-values $0.3$, $0.1$ and $0.05$ respectively. When comparing to a population of 69 events, we see no statistical evidence for HoMs in GW151226, GW170729 and GW190519_153544.
@@ III.2 Correlation between the precession phase and probability of precession
The precession phase [^2] The precession phase, $\phi_{JL}$, is sometimes denoted in the literature as $\alpha$. We denote the $2\pi$-periodic version of the precession phase by $\hat{\phi}_{JL}$ to match the notation used in Ref. [94]. , $\phi_{JL}$, tracks the azimuthal angle of the orbital angular momentum $\mathbf{L}$ around the total angular momentum $\mathbf{J}$, and is typically defined at a reference frequency $f=20\,\mathrm{Hz}$ [^3] In this work we use the same settings as the original LVK analyses. This means that for most of the confident binary black hole mergers considered, we use a reference frequency of $f=20\,\mathrm{Hz}$. However, 22 candidates used a lower reference frequency: $f\in[3,5,10,16]\,\mathrm{Hz}$. For details, see the GWTC-3 data release [4]. When precession is not observable, the precession phase cannot be constrained and Bayesian analyses will return an uninformative measurement of $\phi_{JL}$. However, if the GW has observable precession and no HoMs, Equation 1 reduces to $h\approx h^{\prime}+h_{p}$, and the precession phase can be measured independently from the orbital phase. Since the orbital phase remains the same for both waveform components, while the contribution from the precession phase differs, constraints can be placed on the value of $\phi_{JL}$.
Recently Ref. [94] found that the precession phase, or rather a $2\pi$-periodic version of the precession phase $\hat{\phi}_{JL}$ [^4] Although the precession phase is typically defined at a certain reference frequency, the $2\pi$-periodic version of the precession phase, introduced in Ref. [94], is independent of the frequency specified since it considers the difference between the phase and the circular mean of the phase., is recovered surprisingly well for several candidates, most notably GW200129, GW190412 and GW151226. In Figure 2, we reproduce the analysis of Ref. [94] but look for correlations between the precession p-value and the precision of the measurement of $\hat{\phi}_{JL}$.
As expected we see that, in general, a smaller precession p-value correlates with a tighter constraint on $\hat{\phi}_{JL}$. GW200129, GW190412 and GW151226 have the tightest constraints on $\hat{\phi}_{JL}$, and they have small precession p-values: $4\times 10^{-5}$, $0.04$ and $0.12$. We see one noticeable exception, GW191109_010717. This candidate has a p-value of 0.018 with a relatively poorly constrained $\hat{\phi}_{JL}$.
@@ III.3 Comparing observations with expectations from population estimates
We now turn to comparing the fraction of observed events that display signatures of precession or HoMs with expectations from the population model described below. We check for consistency by sampling binaries from the population model, calculating the expected distributions of p-values for precession and HoMs, and comparing these distributions with observations. In this work we assume a Powerlaw+peak mass model [95], a Default spin model [96] and a Powerlaw redshift model [97], all described in Appendix B of Ref. [15]. We randomly draw 1000 samples from the posterior distribution on the hierarchical model parameters obtained in Ref. [15]; in effect we simulate 1000 different Universes, each consistent with the data. We then randomly sample merging binaries from each Universe. We hereafter refer to this population model as the LVK population model.
@@
Not all binaries drawn from population models are observable with current template-based search pipelines adopted by the LVK [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110]. Current template-based search pipelines adopted by the LVK use templates which restrict spins to be aligned with the orbital angular momentum and neglect the presence of HoMs (although see Refs.[87, 111, 112, 113, 12] for possible solutions for template-based search pipelines, Refs. [114, 115, 116] for search pipelines that do not use templates, and Ref. [117] which demonstrates that non-template based search pipelines may be more sensitive to precessing binaries than current template-based search pipelines). This means that current template-based searches are not equally sensitive to all binaries, and binaries with significant evidence for precession and HoMs may remain undetected.
@@
As discussed in Section II, multiple Bayesian inference analyses are needed to calculate p-values for precession and HoMs for a single randomly chosen binary. Since performing Bayesian inference is a computationally expensive task, let alone analysing each binary with multiple models, we only analyse randomly chosen binaries with $\rho_{\mathrm{p}}>1$ and $\rho_{33}>1$. Although we expect similarities between populations with $\rho_{\mathrm{p}}>1$ and $\rho_{33}>1$, for example asymmetric component masses [42, 50], we treat them as separate populations in this work. This means that we separately draw two populations: one where binaries have $\rho_{\mathrm{p}}>1$ and another where binaries have $\rho_{33}>1$. We find that $11.5\%$ and $5.6\%$ of the detectable binaries sampled from the LVK population model have $\rho_{\mathrm{p}}>1$ and $\rho_{33}>1$, respectively. To ensure a consistent number of binaries in each population, we initially draw $100$ detectable binaries from each Universe in the LVK population model, giving a total of $10^{5}$ randomly chosen binaries, and downsample to 1000 and 2000 binaries for the precession and HoM analyses respectively. We therefore estimate the expected distribution of p-values for precession and HoMs from population models based on $\sim 120$ simulated GW signals. We take care to renormalize the p-value distributions, based on the expected number of binaries with $\rho_{\mathrm{p}}>1$ and $\rho_{33}>1$.
@@
Figure 3 compares the cumulative distribution of p-values for precession and HoMs with random draws from population estimates. The current LVK population model excludes GW190814 [15], so we also excluded this event from the cumulative distribution. We see that the LVK population model accurately predicts the number of observed events with significant evidence for precession and HoMs, with the distribution of observed p-values lying comfortably within the $1\sigma$ scatter from the population samples. We see large error bars associated with the expectation from population estimates, with the most significant binaries having a large spread of p-values: $10^{-8}<P_{\mathrm{p}}<4\times 10^{-3}$ for precession and $10^{-11}<P_{33}<10^{-2}$ for the $(\ell,|m|)=(3,3)$ multipole, with ranges quoted at the $1\sigma$ level. This is a consequence of analysing a limited number of randomly chosen binaries owing to computational cost. Unsurprisingly, when we include GW190814 in the cumulative distribution, the observed distribution now lies outside of the $1\sigma$ uncertainty. This is expected since current population estimates have vanishingly small probabilities for drawing binaries with extreme mass ratio configurations.
@@
In Figure 4 we show the posterior distribution for the inferred primary spin magnitude, $a_{1}$, and mass ratio, $q=m_{2}/m_{1}\leq 1$, for a binary chosen from the population model. This binary was chosen among several possible candidates ($21$ out of the $\sim 120$ simulated GW signals considered) as it was inferred to have a significant precession p-value: $P_{\mathrm{p}}<4\times 10^{-5}$, and hence a more significant p-value than GW200129. We see that the inferred posterior distribution resembles GW200129, with near extreme spin and $q\sim 0.5$; GW200129 was inferred to have $a_{1}=0.9^{+0.1}_{-0.5}$ and $q=0.6^{+0.4}_{-0.2}$ [43]. Figure 4 also shows the posterior distribution for another randomly chosen binary from the population model; this binary was again chosen among several possible candidates as it was inferred to have a significant HoM p-value: $P_{\mathrm{33}}<2\times 10^{-4}$, and hence a more significant p-value than GW190412. Here, we see that the inferred inclination angle, $\theta_{JN}$, and mass ratio, $q$, resembles GW190412; GW190412 was inferred to have $\theta_{JN}=0.71^{+0.31}_{-0.24}$ and $q=0.28^{+0.12}_{-0.07}$ [38]. In both cases, the true values lie within the inferred 90% credible intervals, but there is a small bias, which is likely a consequence of the specific Gaussian noise realisation chosen. Figure 4 illustrates that GW200129-like and GW190412-like signals are possible in the LVK population model, highlighting that e.g. the observed precession and HoM effects previously reported in GW200129 are consistent with a gravitational-wave signal produced from a quasi-circular merger in the presence of Gaussian noise – non-Gaussian noise artefacts may not be needed to explain the observed posterior distributions [45, 48].
@@
Finally, we can use the distribution of p-values obtained from random draws from the LVK population model to calculate the probability of observing binaries with significant evidence for precession or HoMs. We find that there is a 1 in 50 chance that a detectable binary from the LVK population model will have evidence for precession that is at least as significant as for GW200129, and a 1 in 70 chance that a detectable binary will have evidence of HoMs that is at least as significant as for GW190412. Both are consistent with the observation of 1 binary in 68 (removing GW190814).
@@
In this work, we chose to only compare our observations with one population model. However, we note that since we have only observed a single event with significant evidence for precession, and two events with significant evidence for HoMs, any population model that includes a subpopulation of binaries with a $\sim 2\%$ chance of detectable precession and HoMs will likely be consistent with our observations. This includes the LVK population model, but this is not the only possible model consistent with the data, and care should be taken in making strong statements about the population-wide distribution of spin-orbital misalignments and mass ratios based on a handful of events.
@@ IV Discussion
In this work we investigated whether binary black holes observed by the LVK show evidence for precession and HoMs. We took advantage of the harmonic decomposition of a GW [70] and directly calculated the probability that the observed SNR in precession and HoMs can be explained by Gaussian noise alone. Although the SNRs in precession and HoMs have been used to help explain the physics of several GW candidates [38, 51], they are generally underutilised in population analyses.
@@
We found that only one event shows substantial evidence for precession: GW200129. Meanwhile, two events show substantial evidence for HoMs: GW190412 and GW190814. This implies that any evidence for precession and HoMs in other GW signals is consistent with random fluctuations caused by noise. We then compared our observations with expectations from population models to ensure consistency. We found that the current LVK population model [15] accurately predicts the number of observed events with significant evidence for precession and HoMs. We found that a binary with significant evidence for precession will occur once in every $\sim 50$ detections, and a binary with significant evidence for HoMs will occur once in every $\sim 70$ observations if sources are drawn from the LVK population model. Both fractions are consistent with observations, and the latter is consistent with forecasts from Monte Carlo simulations [120]. However, we note that since significant evidence for precession and HoMs has only been confidently observed in three events, any population model that includes a subpopulation of binaries with $\sim 2\%$ chance of detectable precession and HoMs will likely be consistent with the data.
@@
Although we only find one candidate with significant evidence for precession, this does not necessarily mean that most of the observed binary black holes are non-precessing. Rather it means that the majority of the observed binary black holes do not have individually-observable precession. This could indicate a true lack of precession in binaries whose components spins are aligned with the orbital angular momentum; or component spins that are too small in magnitude for precession to be observable despite misalignment; or extrinsic parameters that are unfavourable for the detectability of precession. Consequently, all we can conclusively say about individual events is that we have observed precession in one candidate, GW200129. GW200129 is therefore likely formed through dynamic encounters (although there are cases where the isolated channel can produce binaries with high spins and significant precession [121]). A possibility is that GW200129 was formed through a hierarchical merger, where spin-induced orbital precession is likely; one component is predicted to be highly spinning and misaligned with the orbit. Indeed, signatures of hierarchical mergers may have already been found among GW observations [26, 122, 123]. Other options include triple black hole systems [124], or formation in active galactic nuclei [125]. Given that we cannot make any statements about precession of the other individual GW candidates, we are unable to comment on their formation mechanism. We note that recently Ref. [126] conducted a dedicated search for precessing binaries using data from O3, and did not find any evidence for additional GW candidates, despite an increase in sensitivity to precessing systems. This similarly implies that the majority of the observed binary black holes do not have individually-observable precession. We also highlight that Ref. [127] equivalently found that precession is only individually identified in GW200129 (using another observable).
@@
We only find two candidates with significant evidence for HoMs, both of which have conclusively unequal component masses. Given that GW190412 does not have observable precession, the isolated formation channel is possible: although the isolated formation channel favours comparable component masses [128, 129], binaries with $q\leq 0.5$ are only disfavoured by a factor of $\sim 10$ and therefore unsurprising given the current population [130, 131, 132]. Hierarchical mergers are also possible, although we would then have expected to observe significant evidence for precession. Both isolated binary evolution models [130, 131, 133] and dynamical formation scenarios [134] struggle to reproduce the extreme mass ratio of GW190814. All binaries radiate power in HoMs [^5] Some HoMs (for example all odd $m$ multipoles) are suppressed for non-spinning, exactly equal mass ratio binaries. However, this is not the case for e.g. precessing binaries. , so the lack of observable power in most of the GW candidates is partly a consequence of their extrinsic parameters.
@@
Throughout this work we only considered the observability of precession and HoMs in binary black holes observed with ground-based detectors operating at sensitivities. Previous works have investigated their observability for next generation detectors, such as Einstein Telescope [135] and Cosmic Explorer [136], and found that HoMs in particular will be observable in a larger fraction of detected GW signals [137, 138, 139].
We predict that Einstein Telescope will observe HoMs once in every 20 candidates, while Cosmic Explorer will observe HoMs once in every 10 candidates. Einstein Telescope will observe precession once in 5 candidates, while Cosmic Explorer will observe precession once in every 4 candidates. To obtain these estimates, we calculate how often randomly chosen binaries from the current population models have $\rho_{\mathrm{p}}$ and $\rho_{\ell,m}$ greater than 2.1 for different detector configurations, following the analysis in Ref. [41]. While it is appropriate to use the inferred mass and spin populations from current observations [15], the current observations are not capable of probing the redshift evolution of the merger rate. Thus we draw binaries according to a representative redshift distribution for next generation detectors, such as the distributions proposed in Refs. [140, 141], and use a low frequency cut-off of 5Hz for the SNR calculation. We also assume that all candidates with SNR greater than 8 are detectable [^6] Although this will likely not be a significant issue for binary black holes, this calculation is not taking into consideration the motion of the GW detector during the observed signal due to the Earth’s rotation.. We do not assume a network of next generation detectors, unlike Ref. [137], so these predictions are a conservative lower bound. We see that Cosmic Explorer is more likely to observe HoMs and precession in GW signals than Einstein Telescope. Current population estimates [15] indicate that the majority of black hole binaries are low mass, with $m_{1}\approx 10M_{\odot}$. From Figs 11 and 12 in Ref. [138] we see that both Cosmic Explorer and Einstein Telescope are sensitive to these mergers out to $z\gtrsim 20$, but that Cosmic Explorer has greater sensitivity to HoMs. This is consistent with our estimate that CE will observe HoMs in a greater fraction of events than ET.
@@
The methods presented in this paper straightforwardly and clearly identifies which GW candidates show significant evidence for precession and HoMs. This method can also be used to verify whether population estimates are consistent with observations. We therefore suggest that future population analyses make use of this approach.